# Heterogeneous Directed Hypergraph Neural Network over abstract syntax tree (AST) for Code Classification

### v3

@@ Abstract
Code classification is a difficult issue in program understanding and automatic coding. Due to the elusive syntax and complicated semantics in programs, most existing studies use techniques based on abstract syntax tree (AST) and graph neural network (GNN) to create code representations for code classification. These techniques utilize the structure and semantic information of the code, but they only take into account pairwise associations and neglect the high-order correlations that already exist between nodes in the AST, which may result in the loss of code structural information. On the other hand, while a general hypergraph can encode high-order data correlations, it is homogeneous and undirected which will result in a lack of semantic and structural information such as node types, edge types, and directions between child nodes and parent nodes when modeling AST. In this study, we propose to represent AST as a heterogeneous directed hypergraph (HDHG) and process the graph by heterogeneous directed hypergraph neural network (HDHGN) for code classification. Our method improves code understanding and can represent high-order data correlations beyond paired interactions. We assess heterogeneous directed hypergraph neural network (HDHGN) on public datasets of Python and Java programs. Our method outperforms previous AST-based and GNN-based methods, which demonstrates the capability of our model.
@@
hypergraph, heterogeneous graph, code classification, graph neural networks, program understanding
@@ I Introduction
Although these AST-based and graph-based techniques employ the structural information of source code and demonstrate their effectiveness, there is a problem that they only take into the pairwise relationships and ignore the possible high-order correlations between AST nodes. For example, when code is parsed into an AST, each parent AST node has child AST nodes belonging to various fields or called attributes. A parent node may have several child nodes under the same field, and these nodes have high-order correlations with one another. Fig. 1 depicts an example of a python code snippet. The corresponding AST generated by the official python ast module [^1] https://docs.python.org/3/library/ast.html. is illustrated in Fig. 2(a). As we can see, the “Module” is the root node of the AST. It has two child nodes, “Assign” and “Expr” which belong to the field named “body.” When modeling the correlations between the three nodes, previous approaches only consider the pairwise relationships, i.e., the pair of “Module” and “Assign” and the pair of “Module” and “Expr,” as demonstrated in Fig. 3(a). The high-order data correlation that “Assign” and “Expr” both belong to the “body” of “Module” as shown in Fig. 3(b) is dismissed may result in the loss of code structural information.
@@
To tackle the above problems, we suggest a heterogeneous directed hypergraph (HDHG) to model the AST and a HDHGN for code classification. First, we propose to use heterogeneous directed hyperedge to show the relationship of AST nodes, the example of Fig. 3 is shown in Fig. 4(b). Second, we combine deep learning techniques from hypergraph neural networks and heterogeneous graph neural networks to create the HDHGN, and we also add operations to process directed hyperedge.
We evaluate our method on public datasets Python800 and Java250 [17]. Our model gets 97% in accuracy on Python800 and 96% on Java250 which outperforms previous state-of-the-art AST-based and GNN-based work. Our study demonstrates the utility of HDHG and HDHGN for code classification.
@@
  • We assess our model on public datasets and compare it with previous SOTA AST-based and graph-based methods.
@@ II Related Work
Code classification is to classify codes based on their functions. Different from natural language, code has structural information. As a result, several works adopt AST by various techniques. Mou et al. [1] is one of the first works to suggest a Tree-Based Convolutional Neural Network (TBCNN) in code classification. Alon et al. propose code2seq [7] and code2vec [8] to deconstruct code to a collection of paths in its AST. J. Zhang et al. [5] propose a novel neural called ASTNN for source code representation for code classification and clone detection. N. D. Q. Bui et al. [18] propose a novel method named TreeCaps by fusing capsule networks with TBCNN in code classification.
With the popularity of GNN, more works apply kinds of GNN in code classification based on AST to strengthen the comprehension of code structures. M. Allamanis et al. [13] first construct graphs from source code by adding edges like control flow and data flow to AST and employing a gated graph neural network (GGNN) to process program graphs. V. Hellendoorn et al. [19] propose a model called GREAT based on the transformer architecture by extracting global relational information from code graphs. D. Vagavolu et al. [3] propose an approach that can extract and use program features from multiple code graphs. M. Lu et al. [6] improved GGNN in program classification. T. Long [14] proposes a multi-view graph program representation method that combines both data flow and control flow as multiple views and applies GNN to process. W. Wang et al. [4] propose to leverage heterogeneous graphs to show code based on previous work and adopt heterogeneous GNN to process.
@@ IV Methodology
We first convert the code snippet into an AST and construct an HDHG based on it, then put it into our HDHGN. We combine the vector representations for code categorization once we get the network’s node’s vector representation. The overview of our model is demonstrated in Fig. 5.
@@ IV-B3 Heterogeneous Directed Hypergraph Convolution Layer
Our model updates each node vector in one heterogeneous directed hypergraph convolution (HDHGConv) layer. We refer to the framework of two-stage message passing of hypergraph neural network [15] which has two steps: aggregating messages from nodes to hyperedges and aggregating messages from hyperedges to nodes. Contrarily, we add operations that add heterogeneous information and direction information.
@@
Directed hyperedge $e$ gathers message from their tail nodes and head node by transformer attention mechanism [23]. For each message, the attention score is formulated as (5), where $d_{e}$ is the edge type vector.
@@ IV-C Classification
When we obtain the node hidden vectors $h_{1}^{L},h_{2}^{L},\ldots,h_{|N|}^{L}$ from the last layer, we utilize attention pooling to aggregate the information of each node to obtain vector representation $r$ as (13)(14), where $g\in\mathbb{R}^{C_{2}}$ is a learnable vector.
@@ V-B Baselines
We compare our model with AST-based and GNN-based techniques which acquire the best performance in code classification including TBCNN [1], TreeCaps [18], GGNN [13], GREAT [19] and HPG+HGT [4]. TBCNN used a tree-based convolutional neural network to extract features from AST. A model called TreeCaps combines TBCNN and capsule networks. By adding edges like control flow and data flow to AST, the gated graph neural network (GGNN) processes graphs from source code. GREAT is a model extracting global relational information from code graphs based on the transformer architecture. A technique known as HPG+HGT uses a heterogeneous graph transformer to describe code as a heterogeneous graph. We also trained a GCN [25] and a GIN [26] in an experiment to compare.
@@ V-C Experiment settings
We use a parser from the official Python 3.8 ast library and javalang library [^3] https://pypi.org/project/javalang/ to parse the code snippets into ASTs. The embedding vectors are produced by random initialization and learned via training. Our model’s layer number was set to four. The hidden vector dimension size and embedding vector dimension size were both set to 128. We use a narrow multi-head attention [23] mechanism and set the number of heads to eight. We employed Adam optimizer with the learning rate of $5\times 10^{-5}$ to train our model. We set the dropout rate to 0.2. We optimized the hyper-parameters of other baselines for the validation set’s greatest performance. The models were trained for 100 epochs and we saved the models which perform best in validation set.
@@ V-D Results
We use the performance of the model on the test set as the outcome. We select classification accuracy as the metric. We calculate the mean accuracy and standard deviation after five iterations of the experiment. The results are depicted in Table II. Our HDHGN outperforms other methods in both datasets. In Python800, our HDHGN is 2.88% higher than the best baseline. In Java250, our model outperforms baseline models by at least 2.47%. This demonstrates that our model utilizes the semantic and structural features of code AST more effectively than previous approaches.
@@
AST-based TBCNN 91.16 ±$\pm 0.10$ 90.47 ±$\pm 0.10$
TreeCaps 90.33 ±$\pm 0.11$ 91.38 ±$\pm 0.11$
GNN-based GCN 91.94 ±$\pm 0.10$ 90.01 ±$\pm 0.10$
GIN 93.23 ±$\pm 0.10$ 90.72 ±$\pm 0.10$
GGNN 90.13 ±$\pm 0.11$ 89.97 ±$\pm 0.11$
GREAT 93.33 ±$\pm 0.12$ 93.17 ±$\pm 0.11$
@@
HDHGN 97.87 ±$\pm 0.10$
- hyperedge 94.79 ±$\pm 0.11$
- heterogeneous information 95.23 ±$\pm 0.11$
- direction 95.49 ±$\pm 0.10$
@@ VI Conclusion
In this study, we propose an HDHGN for code classification. To possibly encode high-order data correlations between nodes in AST, we introduce the use of hypergraphs. Due to the general hypergraph being homogeneous and undirected which will result in a lack of semantic and structural information, we propose to represent AST as a heterogeneous directed hypergraph. We create an HDHGN accordingly to utilize high-order data correlation, heterogeneous information and direction information better than previous methods. We test our model using open Python and Java datasets, and we compare the results to the baselines developed using the SOTA AST and GNN. The experiment demonstrates that our HDHGN outperforms the baselines. Further ablation study describes that the HDHGN enhances the performance of code classification.

### v4

@@ Abstract
Code classification is a difficult issue in program understanding and automatic coding. Due to the elusive syntax and complicated semantics in programs, most existing studies use techniques based on abstract syntax tree (AST) and graph neural networks (GNN) to create code representations for code classification. These techniques utilize the structure and semantic information of the code, but they only take into account pairwise associations and neglect the high-order data correlations that already exist between nodes of the same field or called attribute in the AST, which may result in the loss of code structural information. On the other hand, while a general hypergraph can encode high-order data correlations, it is homogeneous and undirected which will result in a lack of semantic and structural information such as node types, edge types, and directions between child nodes and parent nodes when modeling AST. In this study, we propose a heterogeneous directed hypergraph (HDHG) to represent AST and a heterogeneous directed hypergraph neural network (HDHGN) to process the graph for code classification. Our method improves code understanding and can represent high-order data correlations beyond paired interactions. We assess our heterogeneous directed hypergraph neural network (HDHGN) on public datasets of Python and Java programs. Our method outperforms previous AST-based and GNN-based methods, which demonstrates the capability of our model.
@@
hypergraph, heterogeneous graph, code classification, graph neural networks, code representation
@@ I Introduction
Although these AST-based and GNN-based techniques employ the structural information of source code and demonstrate their effectiveness, there is a problem that they only take into the pairwise relationships and ignore the possible high-order correlations between AST nodes of the same field or called attribute. When code is parsed into an AST, each parent AST node has child AST nodes belonging to various fields. A parent node may have several child nodes under the same field, and these nodes have high-order correlations with one another. Fig. 1 depicts an example of a python code snippet. The corresponding AST generated by the official python ast module [^1] https://docs.python.org/3/library/ast.html. is illustrated in Fig. 2(a). As we can see, the “Module” is the root node of the AST. It has two child nodes, “Assign” and “Expr” which belong to the field named “body.” When modeling the correlations between the three nodes, previous approaches only consider the pairwise relationships, i.e., the pair of “Module” and “Assign” and the pair of “Module” and “Expr,” as demonstrated in Fig. 3(a). The high-order correlation that “Assign” and “Expr” both belong to the “body” of “Module” as shown in Fig. 3(b) is dismissed. This may result in the loss of code structural information.
@@
To tackle above problems, we propose a heterogeneous directed hypergraph (HDHG) to model the AST and an HDHGN for code classification. First, we propose to use heterogeneous directed hyperedge to show the relationship of AST nodes, the example of Fig. 3 is shown in Fig. 4(b). Second, we combine deep learning techniques from hypergraph neural networks and heterogeneous graph neural networks to create the HDHGN, and we also add operations to process directed hyperedge.
We evaluate our method on public datasets Python800 and Java250 [17]. Our model gets 97% in accuracy on Python800 and 96% on Java250 which outperforms previous state-of-the-art (SOTA) AST-based and GNN-based work. Our study demonstrates the utility of HDHG and HDHGN for code classification.
@@
  • We assess our model on public datasets and compare it with previous state-of-the-art (SOTA) AST-based and GNN-based methods.
@@ II Related Work
Code classification is to classify codes based on their functions. Different from natural language, code has structural information. As a result, several works adopt AST by various techniques. Mou et al. [1] is one of the first works to suggest a tree-based convolutional neural network (TBCNN) in code classification. Alon et al. propose code2seq [7] and code2vec [8] to deconstruct code to a collection of paths in its AST. J. Zhang et al. [5] propose a novel neural network called ASTNN for source code representation for code classification and clone detection. N. D. Q. Bui et al. [18] propose TreeCaps by fusing capsule networks with tree-based convolutional neural network (TBCNN) in code classification.
With the popularity of GNN, more works apply kinds of GNN in code classification based on AST to strengthen the comprehension of code structures. M. Allamanis et al. [13] first construct graphs from source code by adding edges like control flow and data flow to AST and employing a gated graph neural network (GGNN) to process program graphs. V. Hellendoorn et al. [19] propose a model called GREAT based on the transformer architecture by extracting global relational information from code graphs. D. Vagavolu et al. [3] propose an approach that can extract and use program features from multiple code graphs. M. Lu et al. [6] improve gated graph neural network (GGNN) in program classification. T. Long et al. [14] propose a multi-view graph program representation method that combines both data flow and control flow as multiple views and applies GNN to process. W. Wang et al. [4] propose to leverage heterogeneous graphs to show code based on previous work and adopt heterogeneous GNN to process.
@@ IV Methodology
We first convert the code snippet into an AST and construct an HDHG based on it, then put it into our HDHGN. We aggregate the vector representations for code categorization once we get the network’s node’s vector representation. The overview of our model is demonstrated in Fig. 5.
@@ IV-B3 Heterogeneous Directed Hypergraph Convolution Layer
Our model updates each node vector in heterogeneous directed hypergraph convolution (HDHGConv) layers. We refer to the framework of two-stage message passing of hypergraph neural network [15] which has two steps: aggregating messages from nodes to hyperedges and aggregating messages from hyperedges to nodes. Differently, we add operations that add heterogeneous information and direction information.
@@
Directed hyperedge $e$ gathers messages from its tail nodes and head node by transformer attention mechanism [23]. For each message, the attention score is formulated as (5), where $d_{e}$ is the edge type vector.
@@ IV-C Classification
When we obtain the node hidden vectors $h_{1}^{L},h_{2}^{L},\ldots,h_{|N|}^{L}$ from the last layer, we utilize attention mechanism to aggregate the information of each node to obtain vector representation $r$ as (13)(14), where $g\in\mathbb{R}^{C_{2}}$ is a learnable vector.
@@ V-B Baselines
We compare our model with AST-based and GNN-based techniques which acquire the best performance in code classification including TBCNN [1], TreeCaps [18], GGNN [13], GREAT [19] and HPG+HGT [4]. TBCNN used a tree-based convolutional neural network to extract features from AST. A model called TreeCaps combines TBCNN and capsule networks. By adding edges like control flow and data flow to AST, the GGNN processes graphs from source code. GREAT is a model extracting global relational information from code graphs based on the transformer architecture. A technique known as HPG+HGT uses a heterogeneous graph transformer to describe code as a heterogeneous graph. We also trained a GCN [25] and a GIN [26] in an experiment to compare.
@@ V-C Experiment settings
We use a parser from the official Python 3.8 ast library and javalang library [^3] https://pypi.org/project/javalang/ to parse the code snippets into ASTs. The embedding vectors are produced by random initialization and learned via training. Our model’s layer number was set to four. The hidden vector dimension size and embedding vector dimension size were both set to 128. We use a multi-head attention [23] mechanism and set the number of heads to eight. We employed Adam optimizer with the learning rate of $5\times 10^{-5}$ to train our model. We set the dropout rate to 0.2. We optimized the hyper-parameters of other baselines for the validation set’s greatest performance. The models were trained for 100 epochs and we saved the models which perform best in validation set.
@@ V-D Results
We use the performance of the model on the test set as the outcome. We select classification accuracy as the metric. We calculate the mean accuracy and standard deviation after five times of the experiment. The results are depicted in Table II. Our HDHGN outperforms other methods in both datasets. In Python800, our HDHGN is 2.88% higher than the best baseline. In Java250, our model outperforms baseline models by at least 2.47%. This demonstrates that our model utilizes the semantic and structural features of code AST more effectively than previous approaches.
@@
AST-based TBCNN 91.16 $\pm 0.10$ 90.47 $\pm 0.10$
TreeCaps 90.33 $\pm 0.11$ 91.38 $\pm 0.11$
GNN-based GCN 91.94 $\pm 0.10$ 90.01 $\pm 0.10$
GIN 93.23 $\pm 0.10$ 90.72 $\pm 0.10$
GGNN 90.13 $\pm 0.11$ 89.97 $\pm 0.11$
GREAT 93.33 $\pm 0.12$ 93.17 $\pm 0.11$
@@
HDHGN 97.87 $\pm 0.10$
- hyperedge 94.79 $\pm 0.11$
- heterogeneous information 95.23 $\pm 0.11$
- direction 95.49 $\pm 0.10$
@@ VI Conclusion
In this study, we propose an HDHGN for code classification. To possibly encode high-order data correlations between nodes of the same field or called attribute in AST, we introduce the use of hypergraphs. Due to the general hypergraph being homogeneous and undirected which will result in a lack of semantic and structural information, we propose an HDHG to represent AST. We propose an HDHGN accordingly to utilize high-order data correlation, heterogeneous information and direction information better than previous methods. We test our model using open Python and Java datasets, and we compare the results to the SOTA baselines developed using the AST and GNN. The experiment demonstrates that our HDHGN outperforms the baselines. Further ablation study describes that the HDHGN enhances the performance of code classification.